# CKD: Contrastive Knowledge Distillation from A Sample-wise Perspective

### v1

@@ Abstract
In this paper, we present a simple yet effective contrastive knowledge distillation approach, which can be formulated as a sample-wise alignment problem with intra- and inter-sample constraints. Unlike traditional knowledge distillation methods that concentrate on maximizing feature similarities or preserving class-wise semantic correlations between teacher and student features, our method attempts to recover the “dark knowledge” by aligning sample-wise teacher and student logits. Specifically, our method first minimizes logit differences within the same sample by considering their numerical values, thus preserving intra-sample similarities. Next, we bridge semantic disparities by leveraging dissimilarities across different samples. Note that constraints on intra-sample similarities and inter-sample dissimilarities can be efficiently and effectively reformulated into a contrastive learning framework with newly designed positive and negative pairs. The positive pair consists of the teacher’s and student’s logits derived from an identical sample, while the negative pairs are formed by using logits from different samples. With this formulation, our method benefits from the simplicity and efficiency of contrastive learning through the optimization of InfoNCE, yielding a run-time complexity that is far less than $\mathcal{O}(n^{2})$, where $n$ represents the total number of training samples. Furthermore, our method can eliminate the need for hyperparameter tuning, particularly related to temperature parameters and large batch sizes. We conduct comprehensive experiments on three datasets including CIFAR-100, ImageNet-1K, and MS COCO. Experimental results clearly confirm the effectiveness of the proposed method on both image classification and object detection tasks. Our source codes will be publicly available at https://github.com/wencheng-zhu/CKD.
@@
Knowledge distillation, contrastive learning, intra- and inter-sample constraint, sample-wise alignment
@@ I Introduction
Knowledge distillation (KD) [68, 15, 14] compresses knowledge without significant performance degradation, making it possible to deploy resource-intensive models on lightweight devices [40, 74, 19]. Given the practical benefits, knowledge distillation has garnered much attention over the past decade, and numerous knowledge distillation approaches have been proposed and further introduced into a wide spectrum of tasks [10, 5, 75], such as image classification [39, 16, 71], object detection [87, 82, 61], and segmentation [45, 62, 81]. While considerable improvements have been achieved, existing knowledge distillation methods still struggle to determine what knowledge to learn and how to learn it [66, 2, 53]. Here, we posit that the student model can effectively replicate the output of the teacher model when each sample is appropriately aligned.
Figure 1: Illustration of three typical approaches, including classic KD [25], CRD [70], and our CKD. Classic KD emphasizes feature similarities between teacher and student features, while CRD focuses on class-wise semantic information. In contrast, the proposed CKD conducts similarity alignment and captures sample-wise semantic structures simultaneously.
Hinton et al. [25] propose a seminal work on knowledge distillation, with the goal of reducing the Kullback-Leibler (KL) divergence between teacher and student logits. This influential research has led to a surge in related studies, most of which can be categorized into logit-based and feature-based approaches. Logit-based approaches distill the output logits of teachers [83, 35]. Representative methods in this category are vanilla KD [25], DML [89], DIST [26], DKD [90], etc. Following vanilla KD [25], there has been a lack of research on logit distillation, mainly because of its inferior performance relative to feature distillation. However, logits contain compact information that is easier to distill. Furthermore, logit-based methods can take advantage of low computational and storage requirements [34, 37, 86]. Our approach also employs logit distillation.
Feature-based approaches seek to replicate features extracted from intermediate layers of a teacher model [42, 20]. Typical methods include FitNet [59], OFD [24], ReviewKD [7], etc. These methods have advanced the field of knowledge distillation. However, due to rote memorization, they often suffer from severe overfitting that leads to low training losses but high test losses. Given that, RKD [55] resembles high-order structural relations among features instead of feature similarities. SP [72] preserves pairwise similarities in representation space. Although feature-based approaches effectively capture feature similarity structure, they may somewhat overlook semantic knowledge. CRD [70] innovatively distills structural representations through contrastive learning. To handle the requirement for large batch sizes in contrastive learning, CRD implements a memory buffer to enable the sampling of many negative samples, resulting in less-than-ideal training efficiency, particularly for dense prediction tasks.
To address the aforementioned structure-preserving issues, we present a Contrastive Knowledge Distillation approach, dubbed CKD, which treats knowledge distillation as an intra- and inter-sample alignment problem. Fig. 1 briefly illustrates the main idea of the proposed approach in comparison with two representative approaches. We clearly observe that the proposed approach pulls logits from the same sample together while pushing logits from different samples apart. Specifically, our method assumes that a student model can well mimic a teacher model by distilling sample-wise information. This idea is straightforward and reasonable. On the one hand, previous works [59, 7] have shown tremendous success by maximizing intra-sample feature similarities. Inspired by these methods, we directly minimize logit differences within the same sample. On the other hand, solely relying on an intra-sample similarity constraint can easily lead to overfitting [57, 77]. To this end, we dive into the inter-sample constraints among different samples for additional information. Finally, we integrate these two constraints into the loss function.
Formally, the intra- and inter-sample constraints can be cast as a sample-wise contrastive learning problem. Compared to existing contrastive learning approaches [70, 4] that preserve the class-wise semantic structure, the sample-wise contrastive framework can distill teacher knowledge in an unsupervised way without label information. Moreover, our method precisely aligns teacher and student logits by exploiting both similarities and dissimilarities within each sample rather than each class. From the sample-wise perspective, this leads to a relatively small number of associated negative samples, without resorting to large batch sizes for sampling the whole negative space. Thus, the sample-wise formulation enables our approach to maintain low computational complexity in the absence of class-wise information. Furthermore, our method employs logit distillation over feature distillation for two main reasons. One reason is that logits contain high-level and compact semantic information, which is particularly suitable for contrastive learning. The other reason is that the dimension of logits, crucial for negative sampling, is lower. Notably, our method is model-agnostic and can be applied to feature distillation methods.
To sum up, the contributions of this paper are summarized into three aspects:
@@
    We present a contrastive knowledge distillation approach from the sample-wise perspective. Our method simultaneously conducts intra-sample similarity and inter-sample dissimilarity alignment between teacher and student logits.
@@
    We formulate these two intra-sample and inter-sample constraints as a contrastive learning problem with well-designed positive and negative pairs. By using the proposed contrastive formulation, our method can be efficiently and effectively trained.
@@
    We evaluate the proposed approach on benchmarks for image classification and object detection. Without bells and whistles, our approach improves the accuracy of vanilla KD by more than 1.4% and 0.98% both on the CIFAR-100 and ImageNet-1K datasets.
@@ II-A Knowledge Distillation
Knowledge distillation is a representative model compression technique [65, 49, 80] that enables small models to attain strong performance of large models [1, 52, 48]. As a pioneer, Hinton et al. [25] introduced the concept of knowledge distillation that transfers “dark knowledge” by minimizing the KL divergence between teacher logits and student logits. Over the past few years, a rich line of knowledge distillation methods has been proposed [30, 11, 12]. These knowledge distillation methods come in two flavors: logit distillation and feature distillation [50, 33].
@@ II-A1 Logit Distillation
Earlier logit distillation approaches attempt to enhance generalization performance through regularization and optimization techniques. However, there has been limited research on early-stage logit distillation due to inferior performance. DML [89] improved the generalization abilities of networks by collaboratively training an ensemble of models through mutual learning. Instead, TAKD [51] employed multi-step distillation via intermediate-sized networks, effectively bridging the knowledge transfer gap between teacher and student networks. ICKD-C [44] utilized grid-level inter-channel correlation to capture the intrinsic distribution and diversity of feature representations. GLD [29] employed a local spatial pooling layer for the simultaneous extraction of finely localized knowledge and holistic representations. DIST [26] leveraged both inter-class and intra-class relations to align probabilistic distributions and reinforce semantic similarities, respectively. DKD [90] decoupled logits into a target class and all non-target classes, thereby enhancing both the effectiveness and flexibility of knowledge transfer. Auto-KD [35] firstly introduced an automated search strategy by employing Monte Carlo tree search to optimize knowledge distillation architectures. CTKD [39] modulated the task difficulty levels of a student network by using a dynamic and learnable temperature setting. MLKD [27] also proposed a multi-level logit distillation framework, which distills instance-level, batch-level, and class-level information. LSKD [67] considered temperature as the standard deviation of logit and applied logit standardization before the softmax prediction. Besides, there are several other studies [57, 9] devoted to explaining the basic principle of knowledge distillation. Our method belongs to logit distillation that leverages contrastive information between the logits of the teacher and student models.
Figure 2: The overall architecture of the proposed CKD. The newly designed triplets are introduced for similarity and semantic alignment. For example, we classify all other samples as negative, regardless of their class. In this formulation, training efficiency can be improved by reusing negative samples.
@@ II-A2 Feature Distillation
Feature distillation emphasizes the maximization of similarities between intermediate representations. FitNet [59] proposed a feature-based approach that extracted trainable student representations to predict teacher representations. But, the excessive proximity between samples often results in overfitting. Hence, AT [31] mimicked attention maps of a teacher model to enhance student performance. Instead, RKD [55] transferred the relational structure among data examples by incorporating both distance-wise and angle-wise distillation objectives. IRG [46] distilled not just instance features but also the often-overlooked instance relationships and cross-layer feature transformations. CCKD [56] captured both instance-level information and correlations between multiple instances for knowledge transfer. SP [72] introduced a similarity-preserving loss to maintain the activation pattern similarities of input pairs. OFD [24] combined a new margin ReLU, an optimized distillation feature position, and a selective L2 distance to enhance network compression. CRD [70] was the first contrastive distillation approach that preserves structural representations among features. ReviewKD [7] distilled feature representations from cross-level connection paths to guide the learning of a student model. SimKD [6] reused the teacher’s classifier and introduced an additional projector to align student features. CAT-KD [17] enhances knowledge distillation’s interpretability by distilling class activation maps. This approach empowers the student model to identify discriminative regions, improving its performance. Feature-based approaches usually outperform logit-based approaches at the expense of additional computational and storage overhead.
@@ II-B Contrastive Learning
Contrastive learning stands out as a powerful and popular branch of self-supervised learning, and plenty of efforts have been devoted to contrastive learning [3, 23]. We merely review relevant works concerned with NCE and InfoNCE.
Gutmann et al. [18] firstly proposed the theory of noise-contrastive estimation, in which the model learns to discriminate between samples from the actual data distribution and those from a noise distribution, thereby revealing distinct features of the data. In noise-contrastive estimation, increasing the volume of negative samples can largely reduce reliance on the quality of the noise distribution. Sohn et al. [64] became the first to introduce the multi-class N-pair loss for deep metric learning. This approach eliminates the necessity to recalculate embedding representations in every iteration by allowing the reuse of embedding representations from negative samples. InstDisc [78] incorporated a memory bank mechanism that retains instance features from the preceding update step, facilitating negative sampling through integration with this non-parametric memory bank. CPC [54] popularized the InfoNCE loss for contrastive representation learning, which enhances similarities between a sample and its corresponding positive samples while decreasing similarities with negative samples. Likewise, CMC [69] leveraged a contrastive loss function to optimize the mutual information across different view representations of the same sample. MoCo [21] implemented a moving average of model parameters to update a dynamic dictionary using a queue. SimCLR [8] employed data augmentation to generate varied inputs of the same image and employed a contrastive loss to enhance the similarity between representations with different augmentations. CLIP [58] proposed a large-scale multimodal model pre-trained through the joint training of a text encoder and an image encoder with the InfoNCE loss. Previous methods have witnessed the success of contrastive learning. However, most contrastive learning approaches require substantial batch sizes of negative samples to optimize training. In this paper, we harness the potential of contrastive learning and design sample-wise triples for efficient and effective training.
@@ III-A Notations
Knowledge distillation tasks aim to transfer knowledge from the teacher model $\mathcal{T}$ to the student model $\mathcal{S}$. To be specific, given inputs $\mathcal{X}=[\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n-1}]$, we calculate teacher logits $\boldsymbol{T}=\mathcal{T}\left(\mathcal{X};\theta_{\mathcal{T}}\right)$ and student logits $\boldsymbol{S}=\mathcal{S}\left(\mathcal{X};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{T}=[\boldsymbol{t}_{0},\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{n-1}]\in\mathbb{R}^{c\times n}$ and $\boldsymbol{S}=[\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n-1}]\in\mathbb{R}^{c\times n}$. Here, $\theta_{\mathcal{T}}$ and $\theta_{\mathcal{S}}$ are model parameters. $\mathbf{t}_{i}$ and $\mathbf{s}_{i}$ represent the teacher and student logits for the $i$-th sample in which $i\in\mathcal{D}$, $\mathcal{D}=[1,2,\ldots,n-1]$. $n$ denotes the sample count and $c$ signifies the logit dimension that corresponds to the number of classes. Formally, the objective of knowledge distillation approaches can be written as:
@@
$\mathcal{L}_{\text{Task}}$ represents the standard cross-entropy loss for image classification, and $\alpha$ balances the classification loss $\mathcal{L}_{\text{Task}}$ and the distillation loss $\mathcal{L}_{\text{KD}}$. In what follows, we elaborate on the proposed knowledge distillation loss $\mathcal{L}_{\text{KD}}$.

### v2

@@ Abstract
In this paper, we propose a simple yet effective contrastive knowledge distillation framework that achieves sample-wise logit alignment while preserving semantic consistency. Conventional knowledge distillation approaches exhibit over-reliance on feature similarity per sample, which risks overfitting, and contrastive approaches focus on inter-class discrimination at the expense of intra-sample semantic relationships. Our approach transfers ”dark knowledge” through teacher-student contrastive alignment at the sample level. Specifically, our method first enforces intra-sample alignment by directly minimizing teacher-student logit discrepancies within individual samples. Then, we utilize inter-sample contrasts to preserve semantic dissimilarities across samples. By redefining positive pairs as aligned teacher-student logits from identical samples and negative pairs as cross-sample logit combinations, we reformulate these dual constraints into an InfoNCE loss framework, reducing computational complexity lower than sample squares while eliminating dependencies on temperature parameters and large batch sizes. We conduct comprehensive experiments across three benchmark datasets, including the CIFAR-100, ImageNet-1K, and MS COCO datasets, and experimental results clearly confirm the effectiveness of the proposed method on image classification, object detection, and instance segmentation tasks.
@@
Knowledge distillation, contrastive learning, intra-sample consistency, inter-sample contrast, logit alignment
@@ I Introduction
Knowledge distillation (KD) [70, 15] effectively compresses knowledge without significant performance degradation, making it possible to deploy complex models on resource-limited devices [40, 76, 19]. Owing to its practical benefits, knowledge distillation has garnered much attention over the past decade [10, 39]. Numerous knowledge distillation approaches have been proposed [14, 89] and introduced into a wide spectrum of tasks [5, 77], such as image classification [16, 73], object detection [84, 63], and segmentation [45, 64, 83]. Despite considerable improvements achieved, existing knowledge distillation methods still struggle to determine what knowledge to distill and how to distill it [68, 2, 54]. Here, we posit that the student model can well replicate the outputs of the teacher model when each sample is properly aligned.
Figure 1: Comparison of three knowledge distillation approaches. Classic KD aligns feature similarities between paired teacher-student samples, while CRD leverages contrastive learning to align class-wise semantics. In contrast, the proposed CKD preserves sample-wise similarity for intra-sample alignment while capturing structural semantics for inter-sample contrast.
Hinton et al. [25] introduced seminal work on knowledge distillation that transfers knowledge by minimizing the Kullback-Leibler (KL) divergence between teacher and student logits. Subsequent research has evolved along two main groups, including logit-based and feature-based approaches. Logit-based approaches mimic output teacher logits [85, 35]. Representative methods are vanilla KD [25], DML [91], DIST [26], and DKD [92]. Despite minimal computational and storage requirements [34, 37, 88], these approaches generally underperform feature-based methods.
Feature-based approaches align intermediate layer representations between teacher and student networks [42, 20]. Typical methods include FitNet [61], OFD [24], and ReviewKD [7]. These methods achieve significant performance, however, they often overfit due to rote memorization, achieving a low training loss but poor generalization. Subsequent methods like RKD [56] address this limitation by modeling high-order feature correlations, whereas SP [74] enhances robustness through pairwise similarity preservation in representation space. While these techniques alleviate over-fitting to some extent, their core focus on structural relationships may neglect the transfer of semantic knowledge. Contrastive learning methods, such as CRD [72], enhance structural representation distillation. However, this approach encounters practical limitations: the reliance on large negative sample batches necessitates a large memory buffer, incurring high computational costs, especially for dense prediction tasks.
To address structure-preserving limitations, we propose a Contrastive Knowledge Distillation approach, dubbed CKD, which unifies intra- and inter-sample distillation in a contrastive learning framework. Fig. 1 shows the main idea of the proposed approach compared to two representative approaches. We can observe that our approach aligns same-sample logits while separating cross-sample logits. Our method draws inspiration from the cross-modal alignment paradigm of CLIP, where we take knowledge distillation as a dual-modality alignment task between teacher and student embedding space with the fixed teacher networks. Unlike conventional approaches [61, 7] that enforce feature mimicking, our framework conducts cross-modal alignment through contrastive training, and this paradigm enables the student to capture the teacher’s semantic representations beyond feature consistency while leveraging the efficient sample utilization strategy of CLIP for distillation effectiveness [58, 79]. However, directly applying CLIP is problematic as the teacher model is fixed. Therefore, we propose sample-wise contrastive learning with dual constraints: the intra-sample distillation enhances logit similarities within individual samples, and the inter-sample distillation increases differences across samples.
The intra- and inter-sample alignment constraints can be formulated as a sample-wise contrastive learning problem. Unlike existing contrastive distillation methods [72, 4] that focus on class-wise semantic structures, the sample-wise framework has two key advantages. First, our method directly optimizes the similarity and dissimilarity of logits within individual sample rather than aligning teacher and student models at the class level. Second, by eliminating the dependency on class annotations, the framework supports fully unsupervised distillation. Notably, a key benefit of this sample-wise framework lies in its computational efficiency. Different from class-wise methods requiring large batch sizes to sample the entire negative space, our approach needs a minimal number of negative samples. Additionally, we prioritize logit distillation over feature distillation for two reasons. On one hand, logits provide compact and high-level semantic representations that are suitable for contrastive learning. On the other hand, compared to features, the lower dimensionality of logits significantly simplifies negative sampling. The framework remains model-agnostic and can be seamlessly integrated with feature distillation approaches.
Our contributions can be summarized into three aspects:
@@
    We propose a contrastive knowledge distillation framework with dual constraints, in which our method simultaneously conducts intra-sample and inter-sample distillation between teacher and student logits.
@@
    We cast these dual intra-sample and inter-sample constraints into a sample-wise contrastive formulation through customized positive and negative pairs, achieving effective and computationally efficient training.
@@
    Experiments on benchmark datasets validate the superiority of our method. Without bells and whistles, our method outperforms vanilla KD by 1.4% and 0.98% on the CIFAR-100 and ImageNet-1K datasets, respectively.
@@ II-A Knowledge Distillation
Knowledge distillation is a widely adopted model compression technique [67, 50, 82] that allows compact models to achieve competitive performance with their larger counterparts [1, 53, 49]. As a pioneer, Hinton et al. [25] introduced the concept of knowledge distillation that transfers “dark knowledge” through minimization of the KL divergence between teacher and student logit distributions. Over the past few years, a rich line of knowledge distillation methods has been proposed [30, 11, 12]. These methods come in two flavors: logit distillation and feature distillation [51, 33].
@@ II-A1 Logit Distillation
Early logit distillation approaches mainly focused on enhancing model generalization through regularization and optimization. However, research on early-stage logit distillation has remained limited due to its inferior performance. Initial attempts like DML [91] improved network generalization via mutual learning across model ensembles, while TAKD [52] bridged the teacher-student capacity gap through multi-step distillation with intermediate networks. Subsequent advances introduced feature representation enhancement, distributional alignment, and automated framework design. ICKD-C [44] captured feature diversity through grid-level inter-channel correlations. GLD [29]combined local spatial pooling for fine-grained knowledge extraction. DIST [26] aligned inter-class and intra-class probabilistic relations, respectively. DKD [92] decoupled logits into a target class and all non-target classes to improve transfer flexibility. Auto-KD [35] employed a Monte Carlo tree search for architecture optimization. CTKD [39] dynamically adjusted task difficulty via learnable temperature scheduling. MLKD [27] proposed a multi-level logit distillation framework that distills instance-level, batch-level, and class-level information. LSKD [69] took temperature as the standard deviation of logit and applied logit standardization before the softmax prediction. Moreover, there are theoretical studies [58, 9] that further elucidate the principles of distillation. Our method belongs to logit distillation that leverages contrastive teacher-student logits to enhance knowledge transfer.
Figure 2: The overall architecture of the proposed Contrastive Knowledge Distillation framework. The framework incorporates newly designed triplets to optimize intra-sample feature similarity and cross-sample semantics simultaneously. For each input instance, the teacher and student logits form positive pairs, while other samples within the mini-batch serve as negative pairs. This formulation significantly enhances training efficiency through computationally efficient negative sample reuse, effectively addressing the memory constraints typically associated with contrastive learning approaches.
@@ II-A2 Feature Distillation
Feature distillation focuses on aligning intermediate representations of teacher and student models. FitNet [61] pioneered feature distillation by using hint layers to match student features to teacher activations, but its strict alignment often caused overfitting. Subsequent methods adopted different alignment strategies. For example, AT [31] mimicked spatial attention maps of a teacher model to emphasize task-relevant regions. Differently, RKD [56] preserved relational structures between samples by using distance-wise and angle-wise distillation objectives. IRG [46] incorporated both instance relationships and cross-layer feature transformations. CCKD [57] captured multi-instance correlations for knowledge transfer. SP [74] introduced a similarity-preserving loss to maintain the activation patterns of input pairs. Instead, OFD [24] introduced a margin ReLU, an optimized distillation feature position, and a selective L2 distance for efficient network compression. ReviewKD [7] leveraged cross-layer feature aggregation to guide the learning of a student model. SimKD [6] reused the teacher’s classifier and introduced an additional projector to align student features. CAT-KD [17] enhanced the interpretability of knowledge distillation approaches by distilling class activation maps to highlight discriminative regions. While feature-based distillation generally outperforms logit-based methods in accuracy, it incurs significant computational overhead due to intermediate feature storage and alignment.
Existing contrastive knowledge distillation methods typically align teacher-student feature representations by enforcing intra-class similarity and inter-class separation through contrastive learning frameworks. For example, CRD [72] treats teacher features as anchors and leverages large-scale negative samples to push student features away from dissimilar classes. CRCD [94], which optimizes complementary relationships to reduce dependency on negatives. While these approaches demonstrate theoretical promise, there are several limitations, including heavy reliance on large-scale negative sample pairs, compromised scalability due to memory-intensive operations in dense prediction tasks, and insufficient consideration of the fixed teacher model with gradient conflicts. To handle these limitations, we propose a sample-wise contrastive learning framework that simultaneously preserves intra-sample feature similarity and cross-sample semantics with the optimized contrastive triplets, effectively addressing the memory constraints.
@@ II-B Contrastive Learning
Contrastive learning stands out as a powerful and popular branch of self-supervised learning, and plenty of efforts have been devoted to contrastive learning [3, 23]. We merely review relevant works that are concerned with noise-contrastive estimation (NCE) and InfoNCE. Gutmann et al. [18] firstly proposed the theory of noise-contrastive estimation, in which the model learns to discriminate between samples from the actual data distribution and those from a noise distribution, revealing distinct features of the data. In noise-contrastive estimation, increasing the volume of negative samples can largely reduce reliance on the quality of the noise distribution. Sohn et al. [66] extended this framework with the multi-class N-pair loss, enabling efficient reuse of negative embeddings across iterations.
Subsequent works optimized negative sampling and scalability. For example, InstDisc [80] introduced a memory bank to store historical instance features, enabling stable negative sampling without recomputation. CPC [55] popularized the InfoNCE loss by formalizing contrastive learning as maximizing mutual information between positive pairs. Likewise, CMC [71] leveraged a contrastive loss function to optimize the mutual information across different view representations of the same sample. MoCo [21] decoupled batch size from negative sample size via a momentum-updated queue and conducted large-scale negative sampling with limited memory. SimCLR [8] demonstrated that strong data augmentation paired with a nonlinear projection head achieves state-of-the-art results, albeit with high computational demands. CLIP [59] proposed a large-scale multimodal model pre-trained through the joint training of a text encoder and an image encoder with the InfoNCE loss. While these methods highlight the efficacy of contrastive learning, they rely heavily on large batches of negative samples to ensure the effective separation of representations. This dependency creates computational bottlenecks, particularly in dense prediction tasks. In this paper, we adapt contrastive learning to knowledge distillation by rethinking negative sampling efficiency. Instead of relying on massive negative batches, we propose sample-wise triples that leverage intra-class and inter-class relationships. Our approach maintains the structural benefits of contrastive learning while reducing memory overhead, making it scalable to complex tasks like segmentation and detection.
@@ III-A Notations
Knowledge distillation is a technique used to transfer knowledge from a trained teacher model $\mathcal{T}$ to a more compact student model $\mathcal{S}$. Here are some notations, given a batch of input samples $\mathcal{X}=[\boldsymbol{x}_{0},\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{n-1}]$, we calculate teacher logits $\boldsymbol{T}=\mathcal{T}\left(\mathcal{X};\theta_{\mathcal{T}}\right)$ and student logits $\boldsymbol{S}=\mathcal{S}\left(\mathcal{X};\theta_{\mathcal{S}}\right)$, where $\boldsymbol{T}=[\boldsymbol{t}_{0},\boldsymbol{t}_{1},\ldots,\boldsymbol{t}_{n-1}]\in\mathbb{R}^{c\times n}$ and $\boldsymbol{S}=[\boldsymbol{s}_{0},\boldsymbol{s}_{1},\ldots,\boldsymbol{s}_{n-1}]\in\mathbb{R}^{c\times n}$. $\theta_{\mathcal{T}}$ and $\theta_{\mathcal{S}}$ are model parameters. $\mathbf{t}_{i}$ and $\mathbf{s}_{i}$ represent the teacher and student the logits for the $i$-th sample in which $i\in\mathcal{D}$, $\mathcal{D}=[1,2,\ldots,n-1]$. $n$ denotes the sample count, and $c$ is the number of classes that corresponds to the logit dimension. The student model $\mathcal{S}$ learns to mimic the teacher’s output. A typical distillation loss combines task loss and distillation loss. Formally, the objective of knowledge distillation approaches can be written as,
@@
$\mathcal{L}_{\text{Task}}$ is the standard cross-entropy loss for image classification, while $\alpha$ balances the classification loss $\mathcal{L}_{\text{Task}}$ and the distillation loss $\mathcal{L}_{\text{KD}}$. In what follows, we will detail the knowledge distillation loss $\mathcal{L}_{\text{KD}}$.